OpenStudy (aaron96):

A 70% concentration of isopropyl alcohol is often used as a disinfectant for medical equipment such as thermometers and stethoscopes. The initial volume of 90% isopropyl alcohol needed to produce a batch of 70% isopropyl alcohol can be calculated using the following formula: (90%)(initial volume) = (70%)(final volume) How many millimeters of 90% isopropyl alcohol are needed to create 1 liter of 70% isopropyl alcohol? A. 0.00129 mL B. 0.78 mL C. 1.29 mL D. 780 mL

OpenStudy (welshfella):

0.90x = 0.70*1 x = 0.70/90 l = 0.78 l or 780 ml

OpenStudy (aaron96):

thank you welshfella